OpenStudy (anonymous):

solve for x: 1+32/(x-2)(x+2)=8/x-2

jhonyy9 (jhonyy9):

so the common denominator is x2 -4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Hi there. Do you want to learn how to solve this or just get an answer?

OpenStudy (anonymous):

solve

OpenStudy (anonymous):

6

jhonyy9 (jhonyy9):

than 1 multiply by (x2 -4) 32 multiply by nothing 8 multiply by (x+2)

OpenStudy (anonymous):

if you want to get rid of the fractions,multiply the whole equation by (x-2)(x+2)

jhonyy9 (jhonyy9):

than x2 -4 +32 = 8(x+2)

jhonyy9 (jhonyy9):

x2 -4 +32=8x+16 x2 -8x +12=0

jhonyy9 (jhonyy9):

x_1,_2=(8+/- sqrt(64-48))/2 =(8+/- 4)/2 = 2 and 6

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

do you know what the excluded values that make these expressions undefined are?

OpenStudy (anonymous):

2 is not a valid solution because then the divisor is undefined

jhonyy9 (jhonyy9):

yes right x not can being never 2 or -2

OpenStudy (anonymous):

so the only solution is x=6

jhonyy9 (jhonyy9):

thank you ktklown